OpenStudy (anonymous):

FIND THE DIMENSION OF A RECTANGULAR VWHOSE PERIMETER IS 90 M. AND WHOSE LEGNTH EXCEEDS IT BY 3 M.?

OpenStudy (anonymous):

Please don't post with all upper case, it signify yelling.

OpenStudy (anonymous):

Let the length be 'x' metres Therefore its breadth is x+3m By data perimeter =2 (length +Breadth) Therfore substituting we have, 2(x+x+3)=90 2(2x+3)=90 4x+6=90 Therefore 4x=84 x=21 metres Answer= Hence the length is 21 metres and breadth is 21=3=24 metres !!!!